Experts, Can I get answer in 1 hour or less . please

Thank you

Special Order: High-Low Cost Estimation
SafeRide, Inc. produces air bag systems that it sells to North American automobile manufacturers. Although the company has a capacity of 300,000 units per year, it is currently producing at an annual rate of 180,000 units. SafeRide, Inc. has received an order from a German manufacturer to purchase 60,000 units at $7.00 each. Budgeted costs for 180,000 and 240,000 units are as follows:

180,000 Units 240,000 Units
Manufacturing costs
Direct materials $450,000 $600,000
Direct labor 315,000 420,000
Factory overhead 1,215,000 1,260,000
Total 1,980,000 2,280,000
Selling and administrative 765,000 780,000
Total $2,745,000 $3,060,000
Costs per unit
Manufacturing $11.00 $9.50
Selling and administrative 4.25 3.25
Total $15.25 $12.75

Sales to North American manufacturers are priced at $19 per unit, but the sales manager believes the company should aggressively seek the German business even if it results in a loss of $5.75 per unit. She believes obtaining this order would open up several new markets for the company's product. The general manager commented that the company cannot tighten its belt to absorb the $345,000 loss ($5.75 × 60,000) it would incur if the order is accepted.

(a) Calculate the net advantage (disadvantage) of accepting the order from the German business.

Use a negative sign with your answer to indicate a net disadvantage, if applicable.

$Answer

(b) Calculate the net advantage (disadvantage) of accepting the order from the German business, assuming the company is operating at full capacity.

Use a negative sign with your answer to indicate a net disadvantage, if applicable.

$Answer
Answer #1

Solution a:

Computation of Net Advantage (Disadvantage) accepting Special order - Safe ride inc.
Particulars Amount
Revenue from special order (60000*$7) $420,000.00
Incremental costs:
Direct material ($600,000 - $450,000) $150,000.00
Direct labor ($420,000 - $315,000) $105,000.00
Factory overhead ($1,260,000 - $1,215,000) $45,000.00
Selling and administrative expenses ($780,000 - $765,000) $15,000.00
Net Advantage (Disadvantage) on accepting special order $105,000.00

Solution b:

Direct material per unit = $450,000 / 180000 = $2.50 per unit

Direct labor per unit = $315,000 / 180000 = $1.75 per unit

Variable overhead per unit = ($1,260,000 - $1,215,000) / (240000 - 180000) = $0.75 per unit

Variable selling and administrative expenses = ($780,000 - $765,000) / (240000 - 180000) = $0.25 per unit

Variable cost per unit = $2.50 + $1.75 + $0.75 + $0.25 = $5.25 per unit

Regular contribution margin per unit = $19 - $5.25 = $13.75 per unit

Computation of Net Advantage (Disadvantage) accepting Special order - Safe ride inc.
Particulars Amount
Revenue from special order (60000*$7) $420,000.00
Incremental costs:
Direct material ($600,000 - $450,000) $150,000.00
Direct labor ($420,000 - $315,000) $105,000.00
Factory overhead ($1,260,000 - $1,215,000) $45,000.00
Selling and administrative expenses ($780,000 - $765,000) $15,000.00
Loss of regular contribution margin (60000*$13.75) $825,000.00
Net Advantage (Disadvantage) on accepting special order -$720,000.00